# Ivy-Fake: A Unified Explainable Framework and Benchmark for Image and Video AIGC Detection

### v4

@@ Abstract
The rapid development of Artificial Intelligence Generated Content (AIGC) techniques has enabled the creation of high-quality synthetic content, but it also raises significant security concerns. Current detection methods face two major limitations: (1) the lack of multidimensional explainable datasets for generated images and videos. Existing open-source datasets (e.g., WildFake, GenVideo) rely on oversimplified binary annotations, which restrict the explainability and trustworthiness of trained detectors. (2) Prior MLLM-based forgery detectors (e.g., FakeVLM) exhibit insufficiently fine-grained interpretability in their step-by-step reasoning, which hinders reliable localization and explanation. To address these challenges, we introduce Ivy-Fake, the first large-scale multimodal benchmark for explainable AIGC detection. It consists of over 106K richly annotated training samples (images and videos) and 5,000 manually verified evaluation examples, sourced from multiple generative models and real-world datasets through a carefully designed pipeline to ensure both diversity and quality. Furthermore, we propose Ivy-xDetector, a reinforcement learning model based on Group Relative Policy Optimization (GRPO), capable of producing explainable reasoning chains and achieving robust performance across multiple synthetic content detection benchmarks. Extensive experiments demonstrate the superiority of our dataset and confirm the effectiveness of our approach. Notably, our method improves performance on GenImage from 86.88% to 96.32%, surpassing prior state-of-the-art methods by a clear margin.
00footnotetext: $\clubsuit$ Equal Contributions.11footnotetext: 🖂Corresponding author: wenhui.dong@pi3lab.com.22footnotetext: $\star$ The complete author list is provided in the appendix.33footnotetext: Code
@@ 1 Introduction
The rapid development of diffusion-based models has triggered an exponential growth in Artificial Intelligence Generated Content (AIGC), like Sora [sora_videoworldsimulators2024], DALL-E [dalle_3_betker2023improving], Imagen [imagegen_saharia2022photorealistic], and Stable Diffusion [rombach2022high], which have redefined state-of-the-art performance in text-to-image synthesis. However, these advances also raise significant security concerns, including the misuse of DeepFake [Qureshi2024DeepfakeFA], document tampering [Qureshi2024DeepfakeFA], and dataset poisoning [10.5555/3294996.3295110]. The increasing realism of synthetic content blurs the boundary between genuine and fabricated media, posing critical challenges for misinformation control, content provenance verification, and the preservation of public trust.
However, most existing approaches primarily focus on binary authenticity classification [demamba, wildfake_2025, aigvdet_bai2024ai, aide_yan2025a], which limits the human interpretability of model predictions [loki_ye2024loki]. Existing benchmarks also remain inadequate for evaluating explainable AIGC detection. Datasets such as AIGCDetectionBenchmark [aigcdetectionbenchmark_zhong2023patchcraft] and GenVideo [demamba] provide only binary labels, while more recent resources like LOKI [loki_ye2024loki] attempt to incorporate fine-grained anomaly annotations across modalities but are still constrained in scale and diversity. On the other hand, current methods such as AIDE [aide_yan2025a] and Demamba [demamba] are limited to a single modality—either image or video—without exploring unified detection across both domains. Similarly, existing datasets exhibit the same fragmentation: FakeBench [liu2024mmfakebench] emphasizes explainable fake image detection but omits video content, whereas FakeClue [fakeclue_wen2025spot] provides extensive image-level annotations yet lacks integrated video data. This fragmentation leads to substantial gaps in benchmarking and hinders unified advancement in explainable multimodal AIGC detection.
@@
In addition to providing Ivy-Fake, we introduce Ivy-xDetector, a model for detecting AI-generated images and videos and explaining the associated artifacts. As demonstrated by Figure 1, unlike other MLLMs, Ivy-xDetector excels at spotting generative artifacts, spatial in images and both spatial and temporal in videos. By integrating multiple spatial and temporal feature extractors, it detects image-level artifacts and video-level temporal inconsistencies with superior accuracy. Through the incorporation of explainable annotations and carefully curated domain-diverse data, our approach achieves state-of-the-art performance using only a fraction of the data required by existing detectors. While prior image/video classifiers [demamba, aide_yan2025a, aigcdetectionbenchmark_zhong2023patchcraft] often rely on millions (M-level) of samples, our method attains superior performance across both image and video AIGC detection tasks with fewer than 200K samples.
@@
  • We propose Ivy-xDetector, a reinforcement learning–based model by Reinforcement Learning with Verifiable Rewards (RLVR) built upon GRPO, which achieves superior performance on multiple image and video detection benchmarks using fewer than 200K samples—significantly outperforming existing methods that rely on millions of data points.
@@
2 Dataset
2.1 Data Collection
The Ivy-Fake dataset is curated from GenVideo [demamba], LOKI [loki_ye2024loki], FakeClue [fakeclue_wen2025spot], WildFake [wildfake_2025], Kinetics-400 [kay2017kineticshumanactionvideo], and MMFakeBench [liu2024mmfakebench]. Since Ivy-Fake is substantially larger than other datasets, we primarily detail our data processing procedures here.
Image/Video Corruption Filtering.
The integrity of images and videos is critical for MLLM training. We observed that the official GenVideo [demamba] repository contained a large number of unreadable videos. To ensure data quality, we employed the official Qwen2.5-VL [bai2025qwen2] I/O library to filter out corrupted files, discarding images and videos that failed to load.
Image/Video Resolution.
@@ 2.1 Data Collection
Selecting challenging and diverse samples is crucial for model robustness. We extracted visual features from images and video frames using CLIP [Radford2021LearningTV] and computed intra-class similarity scores. Images with excessively high similarity were removed, while videos were left unprocessed. To comprehensively activate and enhance the model’s reasoning capabilities across the task, we propose a high-quality data construction pipeline that integrates explainable detection of AI-generated content.
@@
The Ivy-Fake training set comprises approximately 106K image and video samples collected during the cold-start phase from diverse sources to ensure broad coverage of contemporary generative models. A significant portion was sourced from established public datasets, including GenVideo [demamba], LOKI [loki_ye2024loki], FakeClue [fakeclue_wen2025spot], WildFake [wildfake_2025], Kinetics-400 [kay2017kineticshumanactionvideo], MMFakeBench [liu2024mmfakebench], GenImage [genimage_2023], DiffusionForensics [dire_wang2023dire]. The cold-start phase aimed to enhance the granularity of interpretable content. To further improve the detector’s accuracy, an additional 87,500 image/video samples were stratified and sampled from multiple datasets for RLVR training, including MSRVTT [xu2016msrvtt] (9,000 videos), GenImage [genimage_2023] (32,000 images per category), CnnSpot [cnnspot_wang2020cnn] (19,000 samples), DigiFakeAV [liu2025faceswappingdiffusionbaseddigital] (2,000 samples), GenVideo [demamba] (Fake subset) (9,000 samples), DiffusionForensics [dire_wang2023dire] (4,500 samples), WildFake [wildfake_2025] (12,000 samples), and synthetic data generated by Qwen-Image [wu2025qwenimagetechnicalreport] (634 samples). In total, the dataset used for training comprises 194,513 samples.
@@
To further enrich the dataset and better capture synthetic images prevalent in real-world scenarios, additional sadmples were generated using a diverse set of state-of-the-art image generation models. These models span multiple architectures and training paradigms, ensuring broad coverage of visual styles and generative patterns for robust evaluation. Specifically, extra sources includes samples generated by FLUX.1 [labs2025flux1kontextflowmatching], Lora Flux [labs2025flux1kontextflowmatching] and Qwen-Image [wu2025qwenimagetechnicalreport].
To ensure balanced representation across generative models and avoid potential biases, a stratified sampling strategy was employed during data collection. This approach maintained proportional inclusion of samples from each generator type, thereby preserving the diversity of visual content and facilitating fair and comprehensive evaluation across detection models.
@@ 2.2 Data Annotation
Figure 2: Token Length Distributions and Multi-Dimensional Coverage Across Datasets. Left: Distribution of token lengths across datasets; Right: Coverage of multiple dimensions in explainability datasets, extracted using Qwen3-32B [qwen3]. The Prompt can be seen in appendix.
To generate explainable annotations, we employed the multimodal large language model Gemini 2.5 Pro [genimi_team2023gemini], leveraging a knowledge distillation process to produce structured, interpretable outputs. To ensure consistent and structured output, we adapted the distillation template from DeepSeek-R1-Zero [deepseekr1_guo2025deepseek], which requires the model to first articulate a reasoning process before issuing a final determination.
To enhance annotation accuracy and guide the generation of relevant explanations, we incorporated category-level prior knowledge during annotation. Specifically, Gemini was provided with the ground truth authenticity label (i.e., "real" or "fake") and instructed to explain the rationale behind this classification. The user prompt was standardized as follows: This {file_type} is {label}. Explain the reason. To further ensure the quality and consistency of explainable annotations, we conducted a comprehensive human alignment and verification process, as detailed in Appendix 10.
As illustrated in Table 1, file_type indicates the modality of the input—either "image" or "video"—and {label} represents the ground-truth label, assigned as either "real" or "fake". The distilled explanations were further categorized along two major dimensions to facilitate structured analysis: Spatial Features, which comprises eight sub-dimensions and captures artifacts and inconsistencies observable within individual frames or static images. Temporal Features, which includes four sub-dimensions and describes anomalies associated with motion, temporal coherence, and cross-frame consistency. Since still images inherently lack temporal attributes, this category is exclusively applicable to video annotations. These categories were informed by established taxonomies of visual artifacts in generative content, as detailed in prior research [deng2024survey].
@@ 2.3 Comparison with Existing Datasets
FakeBench [liu2024mmfakebench] - 6 Image 3K 3K
VANE-Bench [bharadwaj2024vanebench] 101 5 Image* 2K 1K
LOKI [loki_ye2024loki] 226.5 $\sim$16 Image+Video 3K 0
FakeClue [fakeclue_wen2025spot] 119.6 26 Image 68K 36K
@@
In contrast, datasets such as FakeClue [fakeclue_wen2025spot] and FakeBench [liu2024mmfakebench] provide only limited annotated samples, thereby restricting their applicability primarily to evaluation scenarios rather than serving as robust resources for model training. Similarly, LOKI [loki_ye2024loki], while offering explainability-focused annotations, is confined to a single modality (images) and evaluates across fewer dimensions. This constraint limits its suitability for comprehensive research on multimodal AIGC detection and explainability.
We identified a long-tail distribution in the distilled dataset, with a small portion of samples exhibiting token lengths concentrated at both extremes. To mitigate the influence of these outliers, we retained only the central 90% of the distribution, filtering out the top and bottom 5% of samples, thereby keeping data with token lengths between 60 and 800. As shown in Figure 2 panel (a), Ivy-Fake exhibits the longest token length distribution, with an average token length of 451.7. Additionally, to conduct a fine-grained analysis of the semantic dimensions embedded in the explanatory language across different interpretability datasets, we employ Qwen3-32B [qwen3] for dimension extraction. To enhance the accuracy of the extraction, we embed supplementary contextual information from each dataset into the model input. For example, in LOKI [loki_ye2024loki], the bounding box coordinates are wrapped with <bbox></bbox> tags, while metadata such as coordinates, titles, and fine-grained descriptions are also included as input to the model. Moreover, we set the Temperature to 0.8 and perform three extraction passes for each sample, retaining only the dimensions consistently identified across all runs. The final aggregated results are visualized in panel (b).
Figure 3: Comparison between Ivy-Fake and FakeVLM [fakeclue_wen2025spot] (NeurIPS 2025). The Ivy-Fake dataset provides richer and more fine-grained interpretability dimensions.
As shown in Figure 3, Ivy-Fake outperforms FakeVLM in terms of interpretability. While FakeVLM only explains superficial factors such as lighting and texture, Ivy-Fake considers multiple semantic dimensions, including hair, skin, and background, leading to a more comprehensive understanding of forged content. The more image and video cases can be found in Appendix 15.
@@ 2.4 Quality Control
To mitigate the impact of hallucinations from Gemini 2.5 Pro during question generation, we sampled 8,000 cases for manual review. Each chain-of-thought (CoT) involving Gemini underwent at least two rounds of human verification. A total of ten users, including senior university students and regular participants, contributed to the verification process, which required approximately 1,000 hours in total. Each case was tested by at least two users to ensure robustness across different modalities of synthetic data. After this rigorous filtering, we retained 5,000 cases as the benchmark test set, consisting of 2,500 image samples and 2,500 video samples, with each modality containing 1,250 real and 1,250 fake instances. The complete dataset distribution is provided in Appendix 11.
3 Methodology

### v5

@@ Abstract
The rapid development of Artificial Intelligence Generated Content (AIGC) techniques has enabled the creation of high-quality synthetic content, but it also raises significant security concerns. Current detection methods face two major limitations: (1) the lack of multidimensional explainable datasets for generated images and videos. Existing open-source datasets (e.g., WildFake, GenVideo) rely on oversimplified binary annotations, which restrict the explainability and trustworthiness of trained detectors. (2) Prior MLLM-based forgery detectors (e.g., FakeVLM) exhibit insufficiently fine-grained interpretability in their step-by-step reasoning, which hinders reliable localization and explanation. To address these challenges, we introduce Ivy-Fake, the first large-scale multimodal benchmark for fake image and video detection. It consists of over 106K richly annotated training samples (images and videos) and 5,000 manually verified evaluation examples, sourced from multiple generative models and real-world datasets through a carefully designed pipeline to ensure both diversity and quality. Furthermore, we propose Ivy-xDetector, a multimodel large language model (MLLM) based on reinforcement fine-tuning (RFT), capable of producing explainable reasoning chains and achieving robust performance across multiple fake image and video detection benchmarks.
[…]
@@ 1 Introduction
The rapid development of diffusion-based models has triggered an exponential growth in Artificial Intelligence Generated Content (AIGC), like Sora (sora_videoworldsimulators2024), DALL-E (dalle_3_betker2023improving), Imagen (imagegen_saharia2022photorealistic), and Stable Diffusion (rombach2022high), which have redefined state-of-the-art performance in text-to-image synthesis. However, these advances also raise significant security concerns, including the misuse of DeepFake (Qureshi2024DeepfakeFA), document tampering (Qureshi2024DeepfakeFA; qu2023doctamper; qu2024miml; qu2025ostf), and dataset poisoning (10.5555/3294996.3295110; ying2026beyond). The increasing realism of synthetic content blurs the boundary between genuine and fabricated media, posing critical challenges for misinformation control, content provenance verification, and the preservation of public trust.
However, most existing approaches primarily focus on binary authenticity classification (demamba; wildfake_2025; aigvdet_bai2024ai; aide_yan2025a), which limits the human interpretability of model predictions (qu2026textshield; qu2024omni; loki_ye2024loki; fakehr1). Existing benchmarks also remain inadequate for evaluating explainable AIGC detection. Datasets such as AIGCDetectionBenchmark (aigcdetectionbenchmark_zhong2023patchcraft) and GenVideo (demamba) provide only binary labels, while more recent resources like LOKI (loki_ye2024loki) attempt to incorporate fine-grained anomaly annotations across modalities but are still constrained in scale and diversity. On the other hand, current methods such as AIDE (aide_yan2025a) and Demamba (demamba) are limited to a single modality—either image or video—without exploring unified detection across both domains. Similarly, existing datasets exhibit the same fragmentation: FakeBench (liu2024mmfakebench) emphasizes explainable fake image detection but omits video content, whereas FakeClue (fakeclue_wen2025spot) provides extensive image-level annotations yet lacks integrated video data. This fragmentation leads to substantial gaps in benchmarking and hinders unified advancement in explainable multimodal AIGC detection.
@@
In addition to providing Ivy-Fake, we introduce Ivy-xDetector, a model for detecting AI-generated images and videos and explaining the associated artifacts. Unlike other MLLMs, Ivy-xDetector excels at spotting generative artifacts, spatial in images and both spatial and temporal in videos. By integrating multiple spatial and temporal feature extractors, it detects image-level artifacts and video-level temporal inconsistencies with superior accuracy. Through the incorporation of explainable annotations and carefully curated domain-diverse data, our approach achieves state-of-the-art performance using only a fraction of the data required by existing detectors. While prior image and video classifiers (demamba; aide_yan2025a; aigcdetectionbenchmark_zhong2023patchcraft) often rely on millions (M-level) of samples, our method attains superior performance across both image and video AIGC detection tasks with fewer than 200K samples.
@@
  • We propose Ivy-xDetector, a reinforcement learning–based model by Reinforcement Learning with Verifiable Rewards (RLVR) built upon Group Relative Policy Optimization (GRPO), which achieves superior performance on multiple image and video detection benchmarks using fewer than 200K samples, significantly outperforming existing methods that rely on millions of data points.
@@
Figure 2: Token Length Distributions and Multi-Dimensional Coverage Across Datasets; Left: Distribution of token lengths across datasets; Right: Coverage of multiple dimensions in explainability datasets, extracted using Qwen3-32B (qwen3). The Prompt can be seen in appendix.
2 Related Work
2.1 Methods for Synthetic Content Detection
Due to growing concerns about the misuse of synthetic data (deng2024survey; qu2025ostf; li2025artificial), research on AI-generated content (AIGC) detection has expanded rapidly in recent years (wang2026forgeryvcr; park2025vidguard; shuai2026detectors; tan2026videoveritas; tan2025veritas; yu2026agentfox; huang2025unishield; Zhou2025AIGIHolmesTE). Most existing models for AI-generated images and videos formulate the task as binary classification, simply predicting whether the content is "real" or "fake." Representative examples include CNN-based AIGVDet (aigvdet_bai2024ai), CNNSpot (cnnspot_wang2020cnn), DIRE (dire_wang2023dire), Mirror (liu2026mirror) and AIDE (aide_yan2025a). Meanwhile, several works have explored the application of multimodal large language models (MLLMs) to AIGC detection, including Synartifact (cao2024synartifact) and Bi-LORA (keita2025bi). However, these approaches largely overlook the importance of interpretability in AIGC detection.
Some efforts attempt to introduce interpretability by leveraging spatial annotations (qu2023doctamper; zhu2025mesorch; qu2024miml; qu2024omni), frequency-domain artifact analysis (zhang2023perceptual) or or hybrid attention (xin2026hytrechybridtemporalawareattention; kong2025token). Nevertheless, the resulting explanations are often difficult for humans to comprehend, as they lack clarity in natural language. This limitation is particularly evident in the video domain, where AI-generated content frequently exhibits obvious flaws, e.g., incoherent frame transitions and object inconsistency, that are easily noticed and reasoned about by humans (deng2024survey). FakeClue (fakeclue_wen2025spot) introduces the use of vision-language models (VLMs) to provide interpretability for image-level detection, but it does not offer a unified framework that integrates both images and videos.
2.2 Datasets for Synthetic Content Detection.
Early datasets for synthetic content detection (zhang2025detecting; jiang2025revisiting; ma2025imdl; du2025forensichub; zhu2025does), such as CNNSpot (cnnspot_wang2020cnn), primarily collected fake images generated by GAN-based models (gan1_goodfellow2014generative; gan2_zhu2017unpaired; gan3_brock2018large). However, with the advent of more advanced generative architectures like diffusion models (diff4_hertz2022prompt; diff5_nichol2021glide; ying2026beyond) and their variants, the authenticity of generated content has significantly increased, making it more challenging for detection models to discern. This has spurred the development of newer datasets, including ArtiFact (cao2024synartifact), GenImage (genimage_2023), GenImagePP (zhou2025breaking) and WildFake (wildfake_2025). GenImage (genimage_2023) comprises images from the 1000 ImageNet (imagenet_ILSVRC15) categories, generated by eight state-of-the-art generators such as Stable Diffusion (rombach2022high) and Midjourney. Nevertheless, these datasets predominantly focus on image-based content. More recently, datasets emphasizing interpretability have also been introduced (zhao2026spinebench). FakeClue (fakeclue_wen2025spot) contains a large amount of image data with explainability annotations but lacks video data. LOKI (loki_ye2024loki) offers data across 26 different categories and includes 18,000 distinct questions; however, its volume of multimodal data is relatively small and primarily suited for evaluation rather than comprehensive model training (ji2025zoom). Therefore, a critical gap exists for a unified benchmark encompassing both image and video modalities to rigorously evaluate the performance of contemporary AIGC detectors.
3 Methodology
3.1 Data Collection
The Ivy-Fake dataset is curated from GenVideo (demamba), LOKI (loki_ye2024loki), FakeClue (fakeclue_wen2025spot), WildFake (wildfake_2025), Kinetics-400 (kay2017kineticshumanactionvideo), and MMFakeBench (liu2024mmfakebench). Since Ivy-Fake is substantially larger than other datasets, we primarily detail our data processing procedures here.
Image/Video Corruption Filtering.
The integrity of images and videos is critical for MLLM training. We observed that the official GenVideo (demamba) repository contained a large number of unreadable videos. To ensure data quality, we employed the official Qwen2.5-VL (bai2025qwen2) I/O library to filter out corrupted files, discarding images and videos that failed to load.
Image/Video Resolution.
@@ 3.1 Data Collection
Selecting challenging and diverse samples is crucial for model robustness. We extracted visual features from images and video frames using CLIP (Radford2021LearningTV) and computed intra-class similarity scores. Images with excessively high similarity were removed, while videos were left unprocessed. To comprehensively activate and enhance the model’s reasoning capabilities across the task, we propose a high-quality data construction pipeline that integrates explainable detection of AI-generated content.
@@
The Ivy-Fake training set comprises approximately 106K image and video samples collected during the cold-start phase from diverse sources to ensure broad coverage of contemporary generative models. A significant portion was sourced from established public datasets, including GenVideo (demamba), LOKI (loki_ye2024loki), FakeClue (fakeclue_wen2025spot), WildFake (wildfake_2025), Kinetics-400 (kay2017kineticshumanactionvideo), MMFakeBench (liu2024mmfakebench), GenImage (genimage_2023), DiffusionForensics (dire_wang2023dire). The cold-start phase aimed to enhance the granularity of interpretable content. To further improve the detector’s accuracy, an additional 87,500 image/video samples were stratified and sampled from multiple datasets for RLVR training, including MSRVTT (xu2016msrvtt) (9,000 videos), GenImage (genimage_2023) (32,000 images per category), CnnSpot (cnnspot_wang2020cnn) (19,000 samples), DigiFakeAV (liu2025faceswappingdiffusionbaseddigital) (2,000 samples), GenVideo (demamba) (Fake subset) (9,000 samples), DiffusionForensics (dire_wang2023dire) (4,500 samples), WildFake (wildfake_2025) (12,000 samples), and synthetic data generated by Qwen-Image (wu2025qwenimagetechnicalreport) (634 samples). In total, the dataset used for training comprises 106K samples.
@@
To further enrich the dataset and better capture synthetic images prevalent in real-world scenarios, additional samples were generated using a diverse set of state-of-the-art image generation models. These models span multiple architectures and training paradigms, ensuring broad coverage of visual styles and generative patterns for robust evaluation. Specifically, extra sources includes samples generated by FLUX.1 (labs2025flux1kontextflowmatching), Lora Flux (labs2025flux1kontextflowmatching) and Qwen-Image (wu2025qwenimagetechnicalreport).
@@ 3.2 Data Annotation
As illustrated in Table 1, “file_type” indicates the modality of the input—either “image” or “video”, and label represents the ground-truth label, assigned as either “real” or “fake”. The distilled explanations were further categorized along two major dimensions to facilitate structured analysis: Spatial Features, which comprises eight sub-dimensions and captures artifacts and inconsistencies observable within individual frames or static images. Temporal Features, which includes four sub-dimensions and describes anomalies associated with motion, temporal coherence, and cross-frame consistency. Since still images inherently lack temporal attributes, this category is exclusively applicable to video annotations. These categories were informed by established taxonomies of visual artifacts in generative content, as detailed in prior research (deng2024survey).
@@ 3.3 Comparison with Existing Datasets
FakeBench (liu2024mmfakebench) - 6 Image 3K 3K
VANE-Bench (bharadwaj2024vanebench) 101 5 Image* 2K 1K
LOKI (loki_ye2024loki) 226.5 $\sim$16 Image+Video 3K 0
FakeClue (fakeclue_wen2025spot) 119.6 26 Image 68K 36K
@@
We identified a long-tail distribution in the distilled dataset, with a small portion of samples exhibiting token lengths concentrated at both extremes (xin2026hytrechybridtemporalawareattention). To mitigate the influence of these outliers, we retained only the central 90% of the distribution, filtering out the top and bottom 5% of samples, thereby keeping data with token lengths between 60 and 800. As shown in Figure 2 panel (a), Ivy-Fake exhibits the longest token length distribution, with an average token length of 451.7. Additionally, to conduct a fine-grained analysis of the semantic dimensions embedded in the explanatory language across different interpretability datasets, we employ Qwen3-32B (qwen3) for dimension extraction. To enhance the accuracy of the extraction, we embed supplementary contextual information from each dataset into the model input. For example, in LOKI (loki_ye2024loki), the bounding box coordinates are wrapped with <bbox></bbox> tags, while metadata such as coordinates, titles, and fine-grained descriptions are also included as input to the model. Moreover, we set the Temperature to 0.8 and perform three extraction passes for each sample, retaining only the dimensions consistently identified across all runs. The final aggregated results are visualized in panel (b).
Figure 3: Comparison between Ivy-Fake and FakeVLM (fakeclue_wen2025spot) (NeurIPS 2025).
Ivy-Fake outperforms FakeVLM in terms of interpretability. While FakeVLM only explains superficial factors such as lighting and texture, Ivy-Fake considers multiple semantic dimensions, including hair, skin, and background, leading to a more comprehensive understanding of forged content.
@@ 3.4 Quality Control
To mitigate the impact of hallucinations from Gemini 2.5 Pro during question generation, we sampled 8,000 cases for manual review. Each chain-of-thought (CoT) involving Gemini underwent at least two rounds of human verification. A total of ten users, including senior university students and regular participants, contributed to the verification process, which required approximately 1,000 hours in total. Each case was tested by at least two users to ensure robustness across different modalities of synthetic data. After this rigorous filtering, we retained 5,000 cases as the benchmark test set, consisting of 2,500 image samples and 2,500 video samples, with each modality containing 1,250 real and 1,250 fake instances.
3.5 Method